### v1

@@ 3.3 Connection to classical meta-analysis methods
In this section, we show that traditional one- and two-stage meta-analyses using ID can be viewed as special cases within our framework.
When individual-level data from all studies are accessible, one-stage meta-analysis uses a mixed-effects model to account for treatment–covariate interactions with pooled data
@@
Various modeling choices can be made through each of these parameters. Each parameter can be treated as a fixed effect (i.e., with distinct fixed values for each study), a common effect (e.g., $\tau_{i}=\tau$ for all $i$), or as a random effect (e.g., $\tau_{i}\sim\mathcal{N}(\tau,\sigma^{2}_{\tau})$). In particular, if $\tau_{i}$ is modeled as a fixed effect, the model yields study-specific treatment estimates rather than a single global treatment effect across studies, since each study is assigned its own coefficient. In our analysis, we assume that both $\tau_{i}$ and $\phi_{i}$ are common effects, i.e., $\tau_{i}=\tau$ and $\phi_{i}=\phi$ for all studies. For the covariates, we allow for either fixed or common effects. Let $F$ denote the set of covariates with fixed effects and $C$ the set with common effects, so that $F\cup C=\{1,\ldots,p\}$. Specifically, for $l\in C$, we impose $\beta_{i,l}=\beta_{l}$ and $\gamma_{i,l}=\gamma_{l}$, corresponding to common effects across studies.
@@
Proposition 2 establishes that the implied weights for the one-stage ID meta-analysis estimator can be obtained by solving a quadratic programming problem that minimizes the squared $l_{2}$ norm of the weights and enforces a special form covariate balance, where fixed effects produce within-study balance and common effects result in balance across all studies. Because the weights are not required to be non-negative, extrapolation is possible. Analogously, the standard two-stage ID meta-analysis is a special case of our two-stage approach in Algorithm 1 by setting $B^{*}=0$, $\delta=0$, $\psi(w)=w^{2}$ and $c_{i}=\widehat{\sigma}_{i}^{2}+\sigma_{\tau}^{2}$, and omitting non-negativity constraints.
The proposed weigthing framework encompasses estimators from both one- and two-stage ID meta-analysis, providing new insights into these classical methods. However, the formulation of Program (12) does not permit to balance covariates toward a personalized target profile. In practice, the study and target populations often differ due to a range of practical reasons. To address some of these differences, it is important to integrate the target covariate profile into the optimization framework. Furthermore, the weights produced by (12) can be negative, which means the final estimator may extrapolate beyond the convex hull of the study samples and is not necessarily sample-bounded. In particular, negative weights can produce sign-reversal phenomena, where a higher outcome, say, for a treated unit can result in a lower outcome average for the treated group, making interpretation difficult in practice.
@@ 4.1 Selection mechanism introduced by sample-boundedness
The subset $R$ is determined in a data-driven manner, and its explicit form is provided in Theorem S.2 in the Supplementary Materials. The geometry of the boundary defining $R$ is governed by the choice of basis functions $B_{k}(\cdot)$. Empirically, units included in $R$ tend to cluster near the target profile $B^{*}$, while those excluded are typically farther from $B^{*}$. Furthermore, Theorem 4 in Section 5 shows that, under correctly specified models, the proposed method can asymptotically detect relevant study donors.
@@ 4.2 Consistency and asymptotic normality
Theorem 2 establishes that $\widehat{\tau}^{*+}_{\mathrm{ID}}$ is consistent under multiple sets of conditions. Analogous to doubly robust estimators, consistency is achieved if the model for the probability weighting factor, the model for the potential outcomes, or an appropriate combination of the two is correctly specified. Likewise, when only AD are available, the proposed estimator $\widehat{\tau}^{*+}_{\mathrm{AD}}$ exhibits similar multiple consistency conditions, as detailed in Section A.2 of the Supplementary Materials.
@@ 5 Diagnostics for study influence and target support
In particular, determining which studies serve as donors and quantifying their influence on the final estimate is crucial. To aid this assessment, we recommend plotting the sample weights for each observation against their indices and covariate values. These simple graphical diagnostics offer clear visualizations of donor contributions and help evaluate how well the contributing studies align with the target population.
Theorem 1 demonstrates that imposing non-negativity constraints on the weights implicitly induces a selection mechanism, restricting the analysis to a subset of units. Building on this result, Theorem 4 further shows that, under correct specification of the probability weighting models, the proposed approach actually selects units in the support of the target population, while systematically excluding those outside it.
@@ 6 Simulation study
In this section, we assess the finite sample performance of our proposed method in comparison with other related estimators using simulated data. We begin by examining a setting where the support of the target population fully overlaps with that of the study population. The simulation design is adapted from Dahabreh et al. (2023) and includes a total of $n+n^{*}=10,000$ individuals. For individuals in the study population, we observe the source study indicator, treatment assignment, three baseline covariates, and a continuous outcome. Full details of the data generation process are provided in Section A.4 of the the Supplementary Materials. For each simulated dataset, we implemented five estimators: (i) the proposed personalized and sample-bounded estimator $\widehat{\tau}^{*+}_{\mathrm{ID}}$, obtained from Program (7); (ii) the personalized and unbounded estimator $\widehat{\tau}^{*-}_{\mathrm{ID}}$, computed from the relaxation of (7) without the non-negativity constraints; (iii) the g-formula estimator $\widehat{\tau}^{\text{g}}$; (iv) the weighting estimator $\widehat{\tau}^{\mathrm{inv}}$, which assigns weights based on the inverse of the product of the estimated propensity score and the estimated probability of selection into the study population; and (v) the augmented estimator $\widehat{\tau}^{\text{aug}}$, which combines the g-formula and weighting approaches. The last three estimators are discussed in Dahabreh et al. (2023). We note that estimation of $\widehat{\tau}^{\mathrm{inv}}$ and $\widehat{\tau}^{\text{aug}}$ requires access to individual-level covariate data from the target population, whereas our estimator $\widehat{\tau}^{*+}_{\mathrm{ID}}$ does not. For the proposed estimators $\widehat{\tau}^{*+}_{\mathrm{ID}}$ and $\widehat{\tau}^{*-}_{\mathrm{ID}}$, we use the $l_{2}$ norm of the weights as the dispersion measure and require exact covariate balance (i.e., $\delta=0$). Here, the unbounded estimator $\widehat{\tau}^{*-}_{\mathrm{ID}}$ is mathematically equivalent to the g-formula estimator $\widehat{\tau}^{\text{g}}$, hence we omit results for $\widehat{\tau}^{\text{g}}$.
@@
Figure 2 compares the weights of $\widehat{\tau}^{*+}_{\mathrm{ID}}$ and $\widehat{\tau}^{*-}_{\mathrm{ID}}$ in the partially overlapping case, using data from a single simulation run. Each point represents an observation, with point size proportional to its assigned weight, color indicating the sign of the weight, and shape denoting whether the observation is within the target support. The target profile is marked by a green asterisk. For the unbounded estimator $\widehat{\tau}^{*-}_{\mathrm{ID}}$, a substantial number of observations outside the target support receive negative weights, reflecting the extrapolative nature of this estimator. In contrast, the sample-bounded estimator $\widehat{\tau}^{*+}_{\mathrm{ID}}$ assigns zero weights to most observations outside the target support, concentrating nearly all nonzero weights on observations within the support region. This illustrates how the proposed personalized and sample-bounded approach can be used to detect and retain reliable study donors that align with the target population. Empirically, the set of observations assigned negative weights by the unbounded method (colored red) substantially overlaps with those assigned zero weights by the bounded method (shown in gray), although the correspondence is not exact. This suggests that negative weights in the unbounded estimator may serve as a warning that the corresponding observations are outliers with respect to the target population.
@@
In the fully overlapping case, all estimators perform well and detailed results are provided in Section A.4 of the Supplementary Materials. Table LABEL:tbl:sim2 compares the estimators in the partially overlapping case under various settings. In this case, the personalized and sample-bounded estimator $\widehat{\tau}^{*+}_{\mathrm{ID}}$ estimator exhibits the smallest bias and RMSE, while the other estimators show non-negligible bias, particularly when sample size is small. This bias arises because these methods allow all individuals, including those outside the target support who follow a different outcome model, to influence the final estimate. In contrast, our personalized and sample-bounded estimator carefully selects study donors and achieves negligible biases. This shows the robustness of $\widehat{\tau}^{*+}_{\mathrm{ID}}$ to overlap violations.
@@ 7.1 ID case study: HALT-C trial
Estimator Treatment group Control group Causal effect
@@
Table 2: Results from the HALT-C trial, targeting the population from the largest research center.
Table 2 summarizes the results of the meta-analyses using the HALT-C trial dataset. Numbers in parentheses represent 95% quantile-based confidence intervals from 10,000 bootstrap samples (the normal confidence intervals based on variance estimates are comparable). All four estimators yield similar point estimates that closely align with the empirical benchmarks estimates derived from the randomized data in the largest center. In this case study, the covariate distributions across research centers are highly comparable, suggesting minimal violations of the overlap assumption. As a result, the sample-bounded estimator performs similarly to the unbounded estimators. This finding reinforces the intuition that when overlap is strong, sample-bounded methods can achieve equivalent efficiency and accuracy while preserving interpretability and robustness against extrapolation. In settings with more severe violations of overlap, we expect the advantages of our sample-bounded approach to become more pronounced.
@@ 7.2 AD case study: psychotherapy for depression
Figure 3 compares these two estimators. The plots illustrate the donor detection and extrapolation behavior of the unbounded method. Here, each point represents a study, with size proportional to its weight and color indicating the sign of the weight. The green asterisk marks the selected target study. Under the unbounded estimator, some studies are assigned negative weights, particularly those located in regions with more sessions and smaller age than the target profile. In contrast, the bounded estimator discards these studies by assigning them zero weight, resulting in a more interpretable and sample-bounded selection of donor studies. The geometry of the detection boundary is influenced by the choice of basis functions. In this case, the detection boundary for contributing studies takes the shape of a hyperball as we include the squared terms of the covariates. Table 3 summarizes the results from the meta-analyses for this dataset.
@@ 8 Concluding remarks
Meta-analysis is a powerful tool for synthesizing evidence across study designs and data sources. In this paper, we present a weighting framework for evidence synthesis that encompasses and extends regression-based meta-analysis to accommodate both individual- and aggregate-level data. This approach emphasizes interpretability and robustness by producing sample bounded estimators that avoid extrapolation, and by clarifying the contribution of each constituent piece of evidence to the final estimate. Additionally, this approach provides a principled means for detecting relevant data from the target population, a result that is of independent interest in regression analysis.
Looking ahead, a promising direction is to extend this framework by leveraging machine learning and generative modeling techniques to incorporate unstructured data sources, such as clinical notes and medical images. Furthermore, in contexts where hidden bias may be present, incorporating sensitivity analysis methods, such as those developed by Mathur and VanderWeele (2020), can further enhance the practical utility and robustness of our approach.

### v2

@@ 3.3 Connection to classical meta-analysis methods
In this section, we show that traditional one- and two-stage meta-analyses using ID can be viewed as special cases within our framework. This connection provides new insights into these classical methods by revealing their implicit weighting schemes.
When ID from all studies are available, a common one-stage approach is to fit a mixed-effects model to the pooled data:
@@
These parameters allow for various modeling choices. Each parameter can be treated as a fixed effect (i.e., with distinct fixed values for each study), a common effect (e.g., $\tau_{i}=\tau$ for all $i$), or as a random effect (e.g., $\tau_{i}\sim\mathcal{N}(\tau,\sigma^{2}_{\tau})$). In particular, if $\tau_{i}$ is modeled as a fixed effect, the model yields study-specific treatment estimates rather than a single global treatment effect across studies, since each study is assigned its own coefficient. In our analysis, we assume that both $\tau_{i}$ and $\phi_{i}$ are common effects, i.e., $\tau_{i}=\tau$ and $\phi_{i}=\phi$ for all studies. For the covariates, we allow for either fixed or common effects. Let $F$ denote the set of covariates with fixed effects and $C$ the set with common effects, so that $F\cup C=\{1,\ldots,p\}$. Specifically, for $l\in C$, we impose $\beta_{i,l}=\beta_{l}$ and $\gamma_{i,l}=\gamma_{l}$, corresponding to common effects across studies.
@@
Proposition 2 builds on Chattopadhyay and Zubizarreta (2023) and establishes that the implied weights for the one-stage ID meta-analysis estimator can be obtained by solving a quadratic programming problem that minimizes the squared $l_{2}$ norm of the weights and enforces a special form covariate balance, where fixed effects produce within-study balance and common effects result in balance across all studies. Because the weights are not required to be non-negative, extrapolation is possible. Analogously, the standard two-stage ID meta-analysis is a special case of our two-stage approach in Algorithm 1 by setting $B^{*}=0$, $\delta=0$, $\psi(w)=w^{2}$ and $c_{i}=\widehat{\sigma}_{i}^{2}+\sigma_{\tau}^{2}$, and omitting non-negativity constraints.
The proposed weighting framework encompasses estimators from both one- and two-stage ID meta-analysis, providing new insights into these classical methods. However, the formulation of Program (12) does not permit to balance covariates toward a personalized target profile. In practice, the study and target populations often differ due to a range of practical reasons. To address some of these differences, it is important to integrate the target covariate profile into the optimization framework. Furthermore, the weights produced by (12) can be negative, which means the final estimator may extrapolate beyond the convex hull of the study samples and is not necessarily sample-bounded. In particular, negative weights can produce sign-reversal phenomena, where a higher outcome, say, for a treated unit can result in a lower outcome average for the treated group, making interpretation difficult in practice.
@@ 4.1 Selection mechanism introduced by sample-boundedness
The subset $R$ is determined in a data-driven manner, and its explicit form is provided in Theorem S.2 in the Supplementary Materials. The geometry of the boundary defining $R$ is governed by the choice of basis functions $B_{k}(\cdot)$. Empirically, units included in $R$ tend to cluster near the target profile $B^{*}$, while those excluded are typically farther from $B^{*}$. Furthermore, Theorem 4 in Section 5 shows that, under correctly specified models, the proposed method can asymptotically detect relevant data points or study donors.
@@ 4.2 Consistency and asymptotic normality
Theorem 2 establishes that $\widehat{\tau}^{*+}_{\mathrm{ID}}$ is consistent under multiple sets of conditions. Analogously to doubly robust estimators, consistency is achieved if the model for the probability weighting factor, the model for the potential outcomes, or an appropriate combination of the two is correctly specified. Likewise, when only AD are available, the proposed estimator $\widehat{\tau}^{*+}_{\mathrm{AD}}$ exhibits similar multiple consistency conditions, as detailed in Section A.2 of the Supplementary Materials.
@@ 5 Diagnostics for study influence and target support
In particular, determining which observations or studies serve as donors and quantifying their influence on the final estimate is crucial. To aid this assessment, we recommend plotting the sample weights for each observation against their indices and covariate values. These simple graphical diagnostics offer clear visualizations of donor contributions and help evaluate how well the contributing studies align with the target population.
Theorem 1 demonstrates that imposing non-negativity constraints on the weights implicitly induces a selection mechanism, restricting the analysis to a subset of units. Building on this result, Theorem 4 establishes an important property of our method. Under a correctly specified weighting model, data points within the support of the target population are assigned non-zero weights, while those outside this support receive zero weight and are thus effectively excluded from the analysis. In other words, the proposed approach actually selects units in the support of the target population, while systematically excluding those outside it.
@@ 6 Simulation study
In this section, we assess the finite sample performance of our proposed method in comparison with other related estimators using simulated data. We begin by examining a setting where the support of the target population fully overlaps with that of the study population. The simulation design is adapted from Dahabreh et al. (2023) and includes a total of $n+n^{*}=10,000$ individuals. For individuals in the study population, we observe the source study ID, treatment assignment, three baseline covariates, and a continuous outcome. Full details of the data generation process are provided in Section A.4 of the the Supplementary Materials.
For each simulated dataset, we implemented five estimators: (i) the proposed personalized and sample-bounded estimator $\widehat{\tau}^{*+}_{\mathrm{ID}}$, obtained from Program (7); (ii) the personalized and unbounded estimator $\widehat{\tau}^{*-}_{\mathrm{ID}}$, computed from the relaxation of (7) without the non-negativity constraints; (iii) the g-formula estimator $\widehat{\tau}^{\text{g}}$; (iv) the weighting estimator $\widehat{\tau}^{\mathrm{inv}}$, which assigns weights based on the inverse of the product of the estimated propensity score and the estimated probability of selection into the study population; and (v) the augmented estimator $\widehat{\tau}^{\text{aug}}$, which combines the g-formula and weighting approaches. The last three estimators are discussed in Dahabreh et al. (2023). We note that estimation of $\widehat{\tau}^{\mathrm{inv}}$ and $\widehat{\tau}^{\text{aug}}$ requires access to individual-level covariate data from the target population, whereas our estimator $\widehat{\tau}^{*+}_{\mathrm{ID}}$ does not. For the proposed estimators $\widehat{\tau}^{*+}_{\mathrm{ID}}$ and $\widehat{\tau}^{*-}_{\mathrm{ID}}$, we use the $l_{2}$ norm of the weights as the dispersion measure and require exact covariate balance (i.e., $\delta=0$). Here, the unbounded estimator $\widehat{\tau}^{*-}_{\mathrm{ID}}$ is mathematically equivalent to the g-formula estimator $\widehat{\tau}^{\text{g}}$, hence we omit results for $\widehat{\tau}^{\text{g}}$.
@@
Figure 2 compares the weights of $\widehat{\tau}^{*+}_{\mathrm{ID}}$ and $\widehat{\tau}^{*-}_{\mathrm{ID}}$ in the partially overlapping case, using data from a single simulation run. Each point represents an observation, with point size proportional to its assigned weight, color indicating the sign of the weight, and shape denoting whether the observation is within the target support. The target profile is marked by a green asterisk. For the unbounded estimator $\widehat{\tau}^{*-}_{\mathrm{ID}}$, a substantial number of observations outside the target support receive negative weights, reflecting the extrapolative nature of this estimator. In contrast, the sample-bounded estimator $\widehat{\tau}^{*+}_{\mathrm{ID}}$ assigns zero weights to most observations outside the target support, concentrating nearly all nonzero weights on observations within the support region. This illustrates how the proposed personalized and sample-bounded approach can be used to detect and retain reliable study donors that align with the target population. Empirically, the set of observations assigned negative weights by the unbounded method (colored red) substantially overlaps with those assigned zero weights by the bounded method (shown in gray), although the correspondence is not exact. This suggests that negative weights in the unbounded estimator may serve as a warning that the corresponding observations are outliers with respect to the target population. In contrast, as guaranteed by Theorem 4, our sample-bounded estimator automatically discards these outliers by assigning them zero weight.
In the fully overlapping case, all estimators perform well and detailed results are provided in Section A.4 of the Supplementary Materials. Table 1 compares the estimators in the partially overlapping case under various settings. In this case, the personalized and sample-bounded estimator $\widehat{\tau}^{*+}_{\mathrm{ID}}$ estimator exhibits the smallest bias and RMSE, while the other estimators show non-negligible bias, particularly when sample size is small. This bias arises because these methods allow all individuals, including those outside the target support who follow a different outcome model, to influence the final estimate. In contrast, our personalized and sample-bounded estimator carefully selects study donors and achieves negligible biases. This shows the robustness of $\widehat{\tau}^{*+}_{\mathrm{ID}}$ to overlap violations.
@@
Results based on 10,000 simulation runs. Under “Balance” “Yes” indicates scenarios where trials have, on average, equal sample sizes, while “No” refers to scenarios with unequal trial sample sizes. In the column labeled “$Z$ varies,” “Yes” denotes scenarios where the treatment assignment mechanism differs across trials.
@@ 7.1 ID case study: HALT-C trial
Table 2 summarizes the results of our meta-analysis on the HALT-C dataset. In this strong-overlap setting — evidenced by the comparable covariate distributions across centers — all four estimators produce similar point estimates that closely align with the experimental benchmark. These findings show that our sample-bounded estimator achieves efficiency and accuracy comparable to unbounded methods when there is good overlap, while retaining its interpretability and robustness to extrapolation. We anticipate the benefits of this robustness would become even more pronounced in settings with poorer overlap.
Estimator Treated mean Control mean Effect estimate
@@
Table 2: Estimated mean outcomes and treatment effects for patients from the largest research center in the HALT-C trial. Values in parentheses indicate 95% quantile-based confidence intervals computed from 10,000 bootstrap samples. Confidence intervals based on normal approximation and variance estimates yield similar results.
Crucially, our proposed methodology offers a key capability that distinguishes it from conventional approaches. Whereas standard meta-analyses are restricted to estimating overall average treatment effects, our method provides the flexibility to target and estimate effects for specific patient populations based solely on summary-level data. To illustrate this feature, Table 3 presents causal effect estimates for three distinct patient profiles (the specific attributes of these patients are provided in the Supplementary Materials). The substantial variation in these estimates highlights the personalized nature of our approach and its potential to move beyond population averages to inform individualized treatment decisions.
Patient Covariate profile Effect estimate
Patient 1 Middle-aged male with metabolic syndrome 28.4-28.4 (42.0,14.1)(-42.0,-14.1)
Patient 2 Young female with mild disease 34.5-34.5 (55.6,14.1)(-55.6,-14.1)
Patient 3 Older male with prior drug use and liver damage 20.9-20.9 (37.6,4.04)(-37.6,-4.04)
Table 3: Estimated mean outcomes and treatment effects for three hypothetical patient populations with distinct medically relevant covariate profiles. Values in parentheses indicate 95% quantile-based confidence intervals computed from 10,000 bootstrap samples. Confidence intervals based on normal approximation and variance estimates yield similar results.
@@ 7.2 AD case study: psychotherapy for depression
Figure 3 illustrates the donor detection and extrapolation behavior of these two estimators. In the sub-figures, each point represents a study, with size proportional to its weight and color indicating the sign of the weight. The green asterisk denotes the selected target study. The unbounded estimator assigns negative weights to several studies, particularly those with a higher number of sessions and a lower average age than the target. In contrast, the bounded estimator discards these studies by assigning them zero weight, resulting in a more interpretable and sample-bounded selection of donor studies. The geometry of the detection boundary is influenced by the choice of basis functions. In this case, the detection boundary for contributing studies takes the shape of a hyperball as we include the squared terms of the covariates. Table 4 complements these visual findings by summarizing the quantitative results of the meta-analyses.
@@ 8 Concluding remarks
Meta-analysis is a powerful tool for synthesizing evidence across study designs and data sources. In this paper, we presented a weighting framework for evidence synthesis that encompasses and extends regression-based meta-analysis to accommodate both individual- and aggregate-level data. This approach emphasizes interpretability and robustness by producing sample bounded estimators that avoid extrapolation, and by clarifying the contribution of each constituent piece of evidence to the final estimate. A key theoretical result, which is of independent interest for regression analysis, establishes that the sample-boundedness constraint provides a principled, data-driven mechanism for detecting observations most relevant to a given target population.
Several avenues for future research remain open. A promising direction is to extend this framework by leveraging machine learning and generative modeling techniques to incorporate unstructured data sources, such as clinical notes and medical images. Furthermore, in contexts where hidden bias may be present, incorporating sensitivity analysis methods, such as those developed by Mathur and VanderWeele (2020), would allow researchers to formally assess the robustness of their conclusions to potential unmeasured confounding, further enhancing the practical utility of our approach.
@@ A.5 Additional details and results from the case studies
Table S.4 below provides the specific attributes of the three hypothetical HALT-C patients.
Table S.4: Covariate profiles for three hypothetical HALT-C patients
Covariate Patient 1 Patient 2 Patient 3
Baseline platelet count 165 220 110
Age 55 38 61
Sex (Female) 0 1 0
Previous use of pegylated interferon 1 1 0
Race (White) 1 0 0
White blood cell count 6.2 5.4 5.1
History of injected recreational drug use 0 0 1
Transfusion history 1 0 0
Body mass index (BMI) 34.2 24.1 28.7
Creatinine level 0.9 0.8 1.1
Ever smoked 1 0 1
Previous use of combination therapy 1 1 0
Diabetes 1 0 0
Serum ferritin (ng/mL) 420 180 600
Splenomegaly (Ultrasound evidence) 0 0 1
Ever consumed alcohol 1 0 1
Hemoglobin (g/dL) 13.1 14.8 12.2
Aspartate aminotransferase (AST, U/L) 85 42 132